# On the relationship between the Wasserstein distance and differences in life expectancy at birth

### v2

@@ 1 Introduction
The aim of this paper is calculating the Wasserstein distance for two life table age-at death distributions and comparing the value to the differences in life expectancy derived from the same life tables. The idea of using the Wasserstein distance to examine the similarity of life table age-at death distributions was introduced by Cilek et al. (2023). Their work, however, has mainly focused on applying the Wasserstein distance to two-dimensional distributions. That is, they compared cause-specific age-at-death distributions on the basis of the Wasserstein distance. This paper concentrates on the one-dimensional case that is shown in equation 3. The following analysis aims to demonstrate the relationship between $W_{1}(P,Q)$ and differences in life expectancy.
2 Methods
2.1 Life expectancy at birth
In demography, life expectancy at birth ($e_{0}$) measures the expected number of life years on the basis of a set of age-specific mortality rates. $e_{0}$ is derived from a period life table - a procedure where age-specific mortality rates are first converted to probabilities of dying and then applied to an hypothetical birth cohort in order to calculate the number of persons alive and deceased in each age interval (see Preston et al. (2001)). Life expectancy at birth can be defined as the mean age at death of the life table population,
$e_{0}=\frac{\int_{0}^{\omega}x\cdot d(x)dx}{\int_{0}^{\omega}d(x)dx},$ (4)
where $d(x)$ denote the age-specific number of persons deceased in the life table and $\omega$ defines the last open-age interval in the life table.
However, a more popular formula is to define $e_{0}$ as the area under the survival curve,
$e_{0}=\int_{0}^{\omega}l(x)dx,$ (5)
where $l(x)$ denotes the life table survivorship function.
Usually, the size of the initial life table birth cohort is set to 100 000 persons. The probability of dying at the oldest age is $1$, i.e., everyone dies in a life table and hence, the sum over all $dx$ equals the initial life table population size. When setting the initial size to $1$, the $d(x)$ function can be seen as a probability density function (PDF). Further, it is well-known in survival analysis that the PDF, CDF, and the survivorship function are directly linked,
$\begin{split}S(x)&=1-\int_{0}^{x}f(u)du\\ &=1-F(x),\end{split}$ (6)
where the CDF, $F(x)$, is obtained by integrating the PDF, $f(x)$, from 0 up to age (or time) $x$. The survivorship function is the complement of the CDF and vice versa,
F(x)=1S(x).F(x)=1-S(x). (7)
2.2 Wasserstein distance and life table age-at-death distributions
Since age-at-death distributions are one-dimensional, we can compute their Wasserstein distance using their CDFs,
$W_{1}(P_{A},P_{B})=\int_{0}^{\omega}|F_{A}(x)-F_{B}(x)|dx,$ (8)
In this case, the $W_{1}$ Wasserstein distance is simply the area between the respective CDFs of population $A$ and $B$. Since the CDF is defined as l(x)=1F(x)l(x)=1-F(x), we have
$|F_{A}(x)-F_{B}(x)|=|l_{A}(x)-l_{B}(x)|,$ (9)
so the area between the CDFs is equal to the area between the survivorship functions,
$\begin{split}W_{1}(P_{A},P_{B})&=\int_{0}^{\omega}|F_{A}(x)-F_{B}(x)|dx\\ &=\int_{0}^{\omega}|l_{A}(x)-l_{B}(x)|dx.\end{split}$ (10)
Here we are using the cost function |xy||x-y| to measure transportation cost. Accordingly, the Wasserstein distance can be interpreted as the expected absolute difference in ages when the two distributions are matched optimally.
2.3 Wasserstein distance and the difference in life expectancy at birth
The difference in life expectancy at birth for populations $A$ and $B$ can be expressed as the net difference of their survivorship functions,
$e_{0,A}-e_{0,B}=\int_{0}^{\omega}(l_{A}(x)-l_{B}(x))dx.$ (11)
In contrast to the $W_{1}$ Wasserstein distance, equation 11 does not reflect absolute differences in survivorship because positive and negative deviations in $l(x)$ can cancel each other out. This is appropriate when the goal is to measure the difference in the expected number of life years. However, in empirical cases where one population has higher survivorship at all ages, $l_{A}(x)\geq l_{B}(x)$ for all $x$, the difference in life expectancy equals the $W_{1}$ Wasserstein distance.
Moreover, equation 10 indicates clearly the range of possible Wasserstein distance values. In case of two identical $l(x)$ functions, Wasserstein distance equals zero. If there is no overlap between the two $l(x)$ functions, Wasserstein distance reaches its maximum value. In a life table with single ages from age zero to age 100, the maximum Wasserstein distance value is therefore 100. The one-dimensional Wasserstein distance value can be interpreted in terms of the average age shift that is required to transform one distribution into another. If the distributions differ mainly because one population dies younger, Wasserstein distance reflects the magnitude of the age shift. Further, if the distributions differ only in shape (one more concentrated vs. one more spread out), Wasserstein distance captures the minimum mass movements required to reconcile these shapes.
@@ 2.4 Using the Wasserstein distance for cause-specific mortality comparisons
Multi-decrement life tables allow the derivation of age-at-death distributions by specific causes of death (Preston et al. 2001). As before, we set the life table radix $l_{0}=1$. Summing $d(x,c)$ over all ages $x$ and causes $c$ therefore equals one, ensuring that the resulting cause-specific age-at-death distribution is a proper probability distribution. The Wasserstein distance for the one-dimensional all-cause distribution can be computed analytically using equation 10. In contrast, the two-dimensional Wasserstein distance does not have a closed-form solution. To obtain the optimal transport plan, we can use the Python Optimal Transport (POT) package, which implements the network simplex algorithm to solve the transport problem.
The two-dimensional Wasserstein distance requires specifying a cost matrix that determines how costly it is to move probability mass between age-cause combinations. As described above, we set the cost of moving mass across ages as the absolute age difference. Choosing the cost of movement between causes of death, however, is less straightforward. Setting the cross-cause cost to zero makes movements between causes free, i.e., the resulting Wasserstein distance would then ignore cause-specific differences and measure only the mismatch in age patterns between two distributions. At the opposite extreme, setting the cost to infinity makes movements between causes impossible. In this case, the algorithm solves the optimal transport problem for each cause separately. Intermediate choices are also possible. For example, setting the cause-movement cost equal to one implies that shifting mass by one year of age has the same cost as reallocating mass across causes. The appropriate scale therefore depends on the substantive interpretation of cause-of-death differences. The choice of the cost for moving mass determines how large Wasserstein distance can be. Mathematically, we can define the cost matrix for the two-dimensional Wasserstein distance as,
@@ 3.1 Sampling from the Human Mortality Database
The database includes period life tables for several countries in various years (sometimes going back to the 1800s). Each period life table is available for women, men, and for both sexes combined. It is computationally demanding to compare differences in $e_{0}$ and $W_{1}$ Wasserstein distance for all possible pairs in the dataset. For this reason, we focus on life tables for both sexes combined and sample 5 000 randomly selected country pairs from the subset. The cross-country comparison refers to one randomly select year, i.e., we do not compare period life tables over time. In addition to $W_{1}$ and differences in $e_{0}$, we calculate the non-overlap index - also called Jaccard index - (Shi et al. 2022) and the Kullback-Leibler divergence for each pair. This allows comparing the $W_{1}$ Wasserstein result to two alternatives measures for distributional similarity. One of the 5 000 sampled pairs is Denmark and Belgium in 1887. For this pair, the difference in $e_{0}$ equals the $W_{1}$ Wasserstein distance ($3.33$). The non-overlap index is $0.14$, while the Kullback-Leibler divergence is $0.02$. The results for all 5 000 pairs is depicted in Figure 3(b) and 4.
@@
In most cases, a large $e_{0}$ difference coincides with a large $W_{1}$ Wasserstein distances. However, we do find cases where the gap in $e_{0}$ is very small but $W_{1}$ Wasserstein suggests large distributional differences. This is observed, for example, when comparing England & Wales with Iceland in 1849. Both countries show a $e_{0}$ value of about 37.3. Yet, the age-at-death distributions differ substantially (see Figure 6(b)). While Iceland shows higher infant mortality as compared to England & Wales, mortality is lower for Iceland at older ages. Accordingly, the net difference in survivorship over age is small but there are still large absolute differences in $l(x)$ (see 6(a)).
@@ 4 Conclusion
When the survivorship functions of two populations do not cross, the $W_{1}$ Wasserstein distance is equal to the difference in $e_{0}$. In such cases, the comparison of two age-at-death distributions is no longer simply a comparison of means, but rather the solution to an optimal transport problem. This perspective provides a novel interpretation of differences in life expectancy at birth.
@@ 5 Acknowledgements
I would like to thank Laura Ann Cilek for inspiring conversations about the Wasserstein distance and for technical discussions on its implementation for comparing life table age-at-death distributions. I am also grateful to Pavel Griogoriev, Sebastian Klüsener, Jiaxin Shi, and Wen Su for helpful comments and suggestions that improved this work.

### v3

@@ 1 Introduction
In the same one-dimensional setting with linear cost |xy||x-y|, the Wasserstein distance can equivalently be expressed using the optimal transport map. In this case, the optimal transport problem uses a Monge solution and the optimal transport map $T:\mathbb{R}\to\mathbb{R}$ is given by applying the quantile transformation,
$T(x)=F_{Q}^{-1}\!\left(F_{P}(x)\right).$ (4)
Using this map, the $W_{1}$ distance can be written as,
$W_{1}(P,Q)=\int_{-\infty}^{+\infty}|x-T(x)|\,dP(x),$ (5)
which represents the expected absolute move of probability mass under the optimal transport from $P$ to $Q$ (Peyré and Cuturi 2019, p. 32) .
To the best of my knowledge, the first scholars who applied the optimal transport theory framework to the field of demography are Oeppen et al. (2021) and Cilek et al. (2023). While Oeppen et al. (2021) discussed optimisation in life-saving models of mortality, Cilek et al. (2023) used the Wasserstein distance to investigate cause-specific mortality differences between border regions.
@@
2 Methods
3 Relationship to differences in life expectancy at birth
Let $d_{A}$ and $d_{B}$ be two life table age-at-death distributions with corresponding survivorship functions, $l_{A}$ and $l_{B}$. Further, assume that in both life tables the radix is set to one, $l_{A}(0)=l_{B}(0)=1$, and have $l_{A}(\omega)=l_{B}(\omega)=0$ because the probability of dying at the last-open age interval, $\omega$, is one.
If the survivorship functions do not crossover,
$l_{A}(x)\geq l_{B}(x)\quad\forall x\in[0,\omega],$ (6)
then the $W_{1}$ Wasserstein distance between $d_{A}$ and $d_{B}$ equals the difference in life expectancy at birth,
$W_{1}(d_{A},d_{B})=e_{0,A}-e_{0,B}.$ (7)
3.1 Proof
With a radix of one, the life table age-at-death distribution can be seen as a probability density function (PDF), and $l(x)$ gives the survival function $S(x)$. Consequently, the respective CDF is given by,
F(x)=1l(x).F(x)=1-l(x). (8)
Life expectancy at birth can be expressed as,
$e_{0}=\int_{0}^{\omega}l(x)dx.$ (9)
Hence,
$e_{0,A}-e_{0,B}=\int_{0}^{\omega}\bigl(l_{A}(x)-l_{B}(x)\bigr)dx.$ (10)
As shown above, in equation 3, the $W_{1}$ Wasserstein distance, is defined as,
$W_{1}(d_{A},d_{B})=\int_{0}^{\omega}\bigl|F_{A}(x)-F_{B}(x)\bigr|dx.$ (11)
Since $F_{A}(x)=1-l_{A}(x)$ and $F_{B}(x)=1-l_{B}(x)$ we have,
$\bigl|F_{A}(x)-F_{B}(x)\bigr|=\bigl|(1-l_{A}(x))-(1-l_{B}(x))\bigr|=\bigl|l_{A}(x)-l_{B}(x)\bigr|.$ (12)
Under the condition that both survivorship functions do not crossover, equation 6, the integrand is non-negative. So, the absolute value can be removed,
$W_{1}(d_{A},d_{B})=\int_{0}^{\omega}\bigl|l_{A}(x)-l_{B}(x)\bigr|\,dx=\int_{0}^{\omega}\bigl(l_{A}(x)-l_{B}(x)\bigr)\,dx.$ (13)
Substituting the right-hand-side of equation 13 into equation 10 yields the difference in life expectancy at birth. $\square$
3.2 Life expectancy at birth as the mean of the age-at-death distribution
Instead of expressing life expectancy at birth as the area under the survivorship function,
$e_{0}=\int_{0}^{\omega}l(x)dx,$ (14)
we can also define $e_{0}$ as the mean of the age-at-death distribution. To see the equivalence, we apply integration by parts to 14,
$\int_{0}^{\omega}l(x)dx$ $=\int_{0}^{\omega}l(x)\cdot 1dx$ (15)
$=\bigl[x\,l(x)\bigr]_{0}^{\omega}-\int_{0}^{\omega}x\,l^{\prime}(x)dx$ (16)
$=0-\int_{0}^{\omega}x\,l^{\prime}(x)dx$ (17)
$=\int_{0}^{\omega}x\,\bigl(-l^{\prime}(x)\bigr)dx.$ (18)
By definition, the age-at-death distribution is,
$d(x)=-l^{\prime}(x).$ (19)
For a life table with a radix of one, $\int_{0}^{\omega}d(x)dx=1$. Hence, life expectancy at birth is given by,
$e_{0}=\int_{0}^{\omega}x\,d(x)dx.$ (20)
More generally, normalization yields,
$e_{0}=\frac{\int_{0}^{\omega}x\,d(x)dx}{\int_{0}^{\omega}d(x)dx}.$ (21)
Thus, life expectancy at birth can be written either as the area under the survivorship function or as the mean of the age-at-death distribution. This implies that the $W_{1}$ Wasserstein distance between two age-at-death distributions equals the difference in their respective means, whenever the corresponding survivorship functions do not crossover,
$W_{1}(d_{A},d_{B})=\bar{d_{A}}-\bar{d_{B}}\quad\text{if}\quad l_{A}(x)\geq l_{B}(x)\quad\forall x\in[0,\omega].$ (22)
3.3 Graphical illustration
The figures 2 and 3 demonstrate the relationship between the $W_{1}$ Wasserstein distance and the difference in $e_{0}$ based on empirical data obtained from the Human Mortality Database (HMD 2025). The first figure depicts a mortality comparison over time (US in 1990 vs. US in 2019). The age-at-death distribution has shifted to older ages as shown in the left panel on figure 2. The mean age at death or $e_{0}$ value increased by 3.55 years (from 75.4 to 78.95 years). The corresponding survivorship function are drawn in the right panel of the figure. There are not crossing over as the survivorship function referring to the year 2019 shows higher proportion of survivors at each age. Therefore, the difference between the two survivorship functions (gray-shaded area) summed over age gives both, the $W_{1}$ Wasserstein distance and the $e_{0}$ difference.
Figure 2: The $W_{1}$ Wasserstein distance equals the difference in life expectancy at birth when the survivorship functions do not crossover, USA in 2019 compared with USA in 1990, both sexes combined
Figure 3 demonstrates a case where both measures lead to different results. While the US and Germany show a very similar mean ages at death in 1990 (75.40 vs. 75.35 years), the age-at-distributions do not look alike. The distribution for the US is characterized by a steeper shape. Further, the density of deaths at very high ages is slightly larger in the German life table. Accordingly, the corresponding survivorship functions indicate higher proportions of survivors at younger ages for Americans but lower proportions of survivors for the very old. In this situation, the $l(x)$ functions do cross and both measures provide different values. The distributional differences are reflected in the $W_{1}$ Wasserstein distance. The difference in $e_{0}$, however, reflects the difference in net survival, i.e., the gap in mean longevity.
Figure 3: The $W_{1}$ Wasserstein distance differs from the difference in life expectancy at birth when the survivorship functions crossover, USA in 2019 compared with Germany in 1990, both sexes combined
@@ 3.4 Using the Wasserstein distance for cause-specific mortality comparisons
As suggested by Cilek et al. (2023), the optimal transport framework can be used to study differences in cause-specific mortality distributions. Multi-decrement life tables allow the derivation of age-at-death distributions by specific causes of death (Preston et al. 2001). As before, we set the life table radix $l_{0}=1$. Summing $d(x,c)$ over all ages $x$ and causes $c$ therefore equals one, ensuring that the resulting cause-specific age-at-death distribution is a proper probability distribution. The Wasserstein distance for the one-dimensional all-cause distribution can be computed analytically using equation 11. In contrast, the two-dimensional Wasserstein distance does not have a closed-form solution. To obtain the optimal transport plan, we can use the Python Optimal Transport (POT) package, which implements the network simplex algorithm to solve the transport problem.
The two-dimensional Wasserstein distance requires specifying a cost matrix that determines how costly it is to move probability mass between age-cause combinations. As described above, we set the cost of moving mass across ages as the absolute age difference. Choosing the cost of movement between causes of death, however, is less straightforward. Setting the cross-cause cost to zero makes movements between causes free, i.e., the resulting Wasserstein distance would then ignore cause-specific differences and measure only the mismatch in age patterns between two distributions. At the opposite extreme, setting the cost to infinity makes movements between causes impossible. In this case, the algorithm solves the optimal transport problem for each cause separately. Intermediate choices are also possible. For example, setting the cause-movement cost equal to one implies that shifting mass by one year of age has the same cost as reallocating mass across causes. The appropriate scale therefore depends on the substantive interpretation of cause-of-death differences. The choice of the cost for moving mass determines how large the Wasserstein distance can be. Mathematically, we can define the cost matrix for the two-dimensional Wasserstein distance as,
@@ 4.1 Sampling from the Human Mortality Database
The database includes period life tables for several countries in various years (sometimes going back to the 1800s). Each period life table is available for women, men, and for both sexes combined. It is computationally demanding to compare differences in $e_{0}$ and $W_{1}$ Wasserstein distance for all possible pairs in the dataset. For this reason, we focus on life tables for both sexes combined and sample 5 000 randomly selected country pairs from the subset. The cross-country comparison refers to one randomly select year, i.e., we do not compare period life tables over time. In addition to $W_{1}$ and differences in $e_{0}$, we calculate the non-overlap index - also called Jaccard index - (Shi et al. 2022) and the Kullback-Leibler divergence for each pair. This allows comparing the $W_{1}$ Wasserstein result to two alternatives measures for distributional similarity. One of the 5 000 sampled pairs is Denmark and Belgium in 1887. For this pair, the difference in $e_{0}$ equals the $W_{1}$ Wasserstein distance ($3.33$). The non-overlap index is $0.14$, while the Kullback-Leibler divergence is $0.02$. The results for all 5 000 pairs is depicted in Figure 5(b) and 6.
@@
In most cases, a large $e_{0}$ difference coincides with a large $W_{1}$ Wasserstein distances. However, we do find cases where the gap in $e_{0}$ is very small but $W_{1}$ Wasserstein suggests large distributional differences. This is observed, for example, when comparing England & Wales with Iceland in 1849. Both countries show a $e_{0}$ value of about 37.3. Yet, the age-at-death distributions differ substantially (see Figure 8(b)). While Iceland shows higher infant mortality as compared to England & Wales, mortality is lower for Iceland at older ages. Accordingly, the net difference in survivorship over age is small but there are still large absolute differences in $l(x)$ (see 8(a)).
@@ 5 Conclusion
When the survivorship functions of two populations do not cross, the $W_{1}$ Wasserstein distance is equal to the difference in $e_{0}$. In such cases, the comparison of two age-at-death distributions is no longer simply a comparison of means, but rather the solution to an optimal transport problem. This perspective provides a novel interpretation of differences in life expectancy at birth. Further, the $W_{1}$ Wasserstein distance might be used to study distributional differences between $e_{0}$ and health expectancy at birth. By definition, the survival curve for healthy survivors is smaller or equal to the conventional survival curve. Accordingly, their $W_{1}$ Wasserstein distance is simply the difference between life expectancy at birth and the health expectancy at birth.
@@ 6 Acknowledgements
A slightly different version of this paper was submitted to Demographic Research Formal Relationship section. The paper was rejected as its added value was deemed very limited but there was consensus about the technical correctness. Maybe someone else finds it interesting or useful. I would like to thank Laura Ann Cilek for inspiring conversations about the Wasserstein distance and for technical discussions on its implementation for comparing life table age-at-death distributions. I am also grateful to Pavel Griogoriev, Sebastian Klüsener, Jiaxin Shi, and Wen Su for helpful comments and suggestions.